OpenStudy (anonymous):

Help! Simplify: 1. (-z)^3(3z^4) 2. 5(x^2z^6)^3(5x^2z^6)^-3 Thank you!:)

jimthompson5910 (jim_thompson5910):

what does (-z)^3 simplify to

OpenStudy (anonymous):

-z? I don't know... :/

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

(-z)^3 = (-1*z)^3 (-z)^3 = (-1)^3*(z)^3 (-z)^3 = -1*z^3 (-z)^3 = -z^3

jimthompson5910 (jim_thompson5910):

(-z)^3(3z^4) turns into -z^3*3z^4

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

add the z's but the negative is still messing me up :/

jimthompson5910 (jim_thompson5910):

think of -z^3*3z^4 as -1z^3*3z^4

jimthompson5910 (jim_thompson5910):

so the coefficients are -1 and 3, they multiply to -1*3 = -3

jimthompson5910 (jim_thompson5910):

z^3 times z^4 = z^(3+4) = z^7 so the final answer is -3z^7

OpenStudy (anonymous):

thank you! :) can you help me with the second one as well?

jimthompson5910 (jim_thompson5910):

how far did you get with it?

OpenStudy (anonymous):

does exponents distribute first? if so, does the ^3 distribute to the 5 as well or just the (x^2z^6)?

jimthompson5910 (jim_thompson5910):

the exponent of 3 only applies to terms inside the parenthesis

jimthompson5910 (jim_thompson5910):

so (x^2z^6)^3 means you cube x^2 and z^6

OpenStudy (anonymous):

I got 1/25 after simplifying

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

thank you for helping me! :) you're the best! :)

jimthompson5910 (jim_thompson5910):

you're welcome